Project

General

Profile

Actions

Hall Task List Spring17

"Stands" unless indicated otherwise. "???" means the task definition wasn't clear to the last Ops/Commissioning meeting

Hall

  • Change over water solenoids for the SS diode absorber trays. Water tank fitted – to connect. Need to connect drain - stands
  • Nikhef Hall probes Melissa/Ajit - still have 2 failed – complete
  • Tracker vacuum. Ed - in progress
  • Tracker turbo pumps. Craig - stands
  • Luminosity monitor sensor 3 Craig/Ed - stands
  • Luminosity monitor sensor 1 Craig/Ed - stands
  • Diffuser iris 4 - re-test with field on 7/2/17. Stands
  • ‘Splash guard’ for Tracker compressor plumbing in trench. Stands superseded
  • Tracker compressor change-over/testing.

Code Updates C&M

  • Ramp commands to SSU M coils – remedial action taken.
  • Alarm handler/state machine. Moved to static config – expected complete end week 13th 24th March. miceecserv reboot on Thursday 6th.
  • Water pump GUI password protected, #1891
  • Network errors messages – in progress
    • Need to specifiy a port number for all CArepeaters - thus need to reboot all IOC-type machines. Need to find good time for miceecserv reboot.
    • Possibly related to the Watchdog failure, which hasn't re-appeared.

Electrical

Spectrometer solenoids
  • Commissioning of Change over water solenoids for the SS diode absorber trays- T. Hartnett – complete, new tank installed, plumbing to be completed prior to re-test.
Focus Coil
  • FC Data logger GUI to be tested and verified – ongoing:
    • have access to the QPS data itself; have a plug-in to get this Labview data into Excel, but need to understand the channel mapping
    • Need a Windows PC running Labview client (rdesktop in) - can put on one of the Hall Probe machines
  • PT100’s on FC in R9 to be fitted – ongoing 0 sensors will be glued leave till summer?
  • FC power supply failure – repair with Sorensen
Hydrogen System
  • Cryo-cooler for LH2 system moved from Vacuum control station 9 rack to H2 control room information to the control system is not available in new location, request to resolve. P. Hanlet to prove. Will be restricted to LH2 Control System
  • Hydrogen level sensing commissioning – P. Warburton to advise is on the dedicated LH2 workplan
Drawings and records (CW to contact DL to clarify items below)
  • Check off power distribution records – J. Webb, A. Gallagher ???
  • Records of testing to be made available - A. Gallagher ???
  • Update all Mice drawings and issue to Mice Wiki page in PDF format – J. Webb, I. Mullacrane to advise ???
MICE Hall
  • 24V supply for draw wire strain gauge required. Installation complete, commissioning and testing to be completed in hall. – A. Gallagher, J. Webb
  • Magnet Interlock alarm is inoperative investigation required – DL engineer ???
Preventative maintenance
  • Spectrometer solenoid Backup Helium Valve: Relay required to repeat valve state to the control system as status of valve is only known local to the valve. - I. Mullacrane
  • Cable/Software check needed to prove which magnet Cold-heads are connected to which Cryo-compressors and that GUI has the correct compressor RS232 communications associated with the correct Cold-heads. - I. Mullacrane, J. Webb
  • Review of FC cable numbers. - J. Webb Link Box labelling has been cleaned up
  • Fit plinth to EMR rack. – A. Gallagher
  • Temperature instrumentation upgrade for Decay Solenoid – I. Mullacrane
  • RR2 to be updated in HevaComp – DL engineer
  • Main 5th AC unit adjacent to mice substation reinstated, drawings require mods to show location and feeder details - J. Webb
  • Stability Run for FC and Spectrometer Solenoid magnets – T. Hartnett, S. Griffiths ???

Detectors

  • New Target control PC (RR1) – in progress
  • Restart R78 target testing – in progress

Updated by Nebrensky, Henry about 6 years ago · 8 revisions